Week of 050718

Open Actions from last week:
  • LEMON sensors for
    • FTS - Gavin WAITING ON FIO
  • Gavin: install last FTS version and test
  • Move all users over to castor2 - James
  • Hourly tests (like SFT) - Simone FIST VERSION DONE
  • Jan/Sophie/Laurence : How to run "central" LEMON sensors LEAVE
  • All: Go through the 2nd level support
  • Ben: check oplapro79 to see if system killed SRM process NOTHING FOUND
  • Olof: castor2 - Try with one node moving DONE
  • bdflush parameters. lxshare220d.
  • James/Ben: check oplapro80 for access for ben.

On shift: Jean-Philippe/Jamie

Monday:

Log: Many interventions.
  • FTS005 died NO_CONTACT - came back up
  • /tmp 92% on sc3-fts - Gavin - FTS dumping core
  • castorgridsc - 1 LSF plugin problem.
    • poor throughput in LSF - restarting LSF helps.

New Actions:

  • Gavin: Investigate core dumping on glite-url-copy
  • Check again the operator levels.

Discussion:

  • hourly tests - like to try with new version of lcg_utils
  • Olof: Wait until this afternoon to move castor1 nodes into WAN pool

Tuesday:

Log:

New Actions:

Wednesday

Log:

Actions:

Thursday

Log:

Actions:

Friday

Log:

Actions:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2005-07-19 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback